A boat, which moves at 18 miles per hour in water without a current, goes 80 miles upstream and 80 miles back again in 9 hours. Find the speed of the current to the nearest tenth.
I am very confused on how to work out this problem...

Answer by stanbon(75887)

A boat, which moves at 18 miles per hour in water without a current, goes 80 miles upstream and 80 miles back again in 9 hours. Find the speed of the current to the nearest tenth.
----
Let current speed be "c" .
------------------------------------------------------
Upstream DATA:
distance = 80 miles ; rate = 18-c mph ; time = d/r = 80/(18-c) hrs.
---------------------------------
Downstream DATA:
distance = 80 miles ; rate = 19+c mph ; time = d/r = 80/(18+c) hrs
----------------------------------------
Equation:
time up + time down = 9 hrs
80/(18-c) + 80/(18+c) = 9
-----
80(18+c) + 80(18-c) = 9(18-c)(18+c)
---------------------------------------------
2*80*18 = 9(324-c^2)
320 = 324-c^2
c^2 = 4
c = 2 mph (speed of the current)
------------------------------------
Cheers,
Stan H.